OpenStudy (anonymous):

2a + 2b = c for b How do I solve for the indicated variable?

OpenStudy (anonymous):

if you have no numbers b would only equal an expression if so then the steps are 2a+2b=c 2b=c-2a b=(c-2a)/2

OpenStudy (anonymous):

2a+2b=c -a -2a c-2a=2b c-2a/2 = 2b/2 1/2c-a=b

OpenStudy (anonymous):

Corrected** 2a+2b=c -2a -2a c-2a=2b c-2a/2 = 2b/2 1/2c-a=b